### v2

# Faster and Better 3D Splatting via Group Training

Abstract
3D Gaussian Splatting (3DGS) has emerged as a powerful technique for novel view synthesis, demonstrating remarkable capability in high-fidelity scene reconstruction through its Gaussian primitive representations. However, the computational overhead induced by the massive number of primitives poses a significant bottleneck to training efficiency. To overcome this challenge, we propose Group Training, a simple yet effective strategy that organizes Gaussian primitives into manageable groups, optimizing training efficiency and improving rendering quality. This approach shows universal compatibility with existing 3DGS frameworks, including vanilla 3DGS and Mip-Splatting, consistently achieving accelerated training while maintaining superior synthesis quality. Extensive experiments reveal that our straightforward Group Training strategy achieves up to 30% faster convergence and improved rendering quality across diverse scenarios.
(a) Reconstruction of the “Stump”[1] scene by 3DGS[12].
(b) Reconstruction of the “Dr. Johnson”[9] scene by Mip-Splatting[30].
Figure 1: Improvements and Illustrations of Applying Group Training in 3DGS[12] and Mip-Splatting[30]. Grouping Training (solid line) achieves a significant increase in reconstruction speed and superior scene quality compared to baseline methods (dashed line). In the “Stump”[1] scene, Grouping Training delivers a 28% faster reconstruction while rendering leaf details with greater clarity. In the “Dr. Johnson”[9] scene, Mip-Splatting with Group Training achieves a 33% reduction in reconstruction time and effectively suppresses the generation of floating artifacts.
1 Introduction
Novel view synthesis (NVS) is a pivotal technology in various domains, such as virtual reality [11], augmented reality [31], and autonomous driving [33]. Recently, 3D Gaussian Splatting (3DGS) [12] achieved significant success in NVS due to its real-time [14, 27, 34], high-quality rendering [2, 19]. This high-quality rendering is made possible by using millions of Gaussians [3], each includes a complex set of attributes, namely, position, size, orientation, opacity, and color. These attributes are optimized via multi-view photometric losses.
$\bullet$ Motivation. However, 3DGS also introduces a potential risk during training, as the exponentially growing Gaussians significantly increase the training burden. Consequently, a primary research focus is the development of methods that enable rapid and cost-effective 3D scene reconstruction without compromising the rendering performance.
During the densification stage, the cyclic pruning of Gaussians with low opacity emerges as a straightforward yet potentially effective strategy [12, 26] for enhancing the training efficiency of 3DGS. Specifically, in each round, 3DGS prunes Gaussians whose opacity falls below a designated threshold $\epsilon_{\alpha}$, which is a vital parameter that facilitates control over the number of generated Gaussians. However, this method has inherent limitations. If $\epsilon_{\alpha}$ is set too conservatively, the resultant acceleration will be minimal while an excessively lenient threshold defeats the quality of the NVS. Thus, achieving a balance between accelerating the training and preserving the rendering quality poses a considerable challenge.
$\bullet$ Intuition and contributions. Inspired by the limitations of the cyclic pruning strategy, we propose to cache a portion of the Gaussians instead of directly pruning Gaussians. This “cache” concept can retain the “important” Gaussians while reducing the total number of them. Therefore, we propose a novel Group Training strategy to accelerate the training of 3DGS. Specifically, (1) a portion of Gaussians is cyclically cached during training. These cached Gaussians are temporarily excluded from scene rendering and model training, which greatly reduces the number of Under-training Gaussians and cuts down on training time. (2) Additionally, through cyclic resampling, the cached Gaussians are reintegrated into the 3DGS model, reducing the impact of direct pruning. (3) A crucial element of Group Training is determining how to sample Gaussians to distinguish between under-training and cached Gaussians. A naive sampling approach would be random sampling. Interestingly, we discovered that the distribution of Gaussian opacity values influences both the quantity of Gaussians generated during densification and the rendering speed of the scene. Therefore, we introduce an Opacity-based Prioritized Sampling method, which effectively reduces the generation of redundant Gaussians and improves the training speed of 3DGS. Specifically, the main technical contributions of our work are as follows:
  1. 1.
    We propose a simple yet highly efficient training framework called Group Training for 3D Gaussian Splatting, which is seamlessly integrated into existing 3DGS frameworks, including both 3DGS and Mip-Splatting.
  2. 2.
    Our framework significantly improves the efficiency of scene reconstruction and the quality of NVS.
Figure 2: PSNR and Time Performance of Pruning [12, 26] and Group Training under Varying Hyperparameters. The reconstructions are performed on the “Train” scene [13] using 3DGS [12] as the baseline. Left: The pruning method exhibits substantial instability in optimizing the trade-off between reconstruction efficiency and quality. The hyperparameter sensitivity (Opacity Threshold) presents significant challenges for optimal parameter tuning. Right: Group Training demonstrates consistent improvements in both reconstruction speed and quality, with robust performance across a wide range of hyperparameter values (Caching Ratio), enabling straightforward parameter optimization.
Our code will be made publicly available upon acceptance of this paper ( but is now available in the supplementary materials for reviewing).
Figure 3: The overall framework of Group Training. Group Training involves periodically dividing all Gaussian primitives. Specifically, at regular iteration intervals, Gaussians from all groups are merged before rendering the training view. Subsequently, all Gaussian primitives are categorized into the Under-training Group and the Caching Group according to a specified sampling strategy. Before the next grouping, the Under-training Group is utilized for gaussian densification (Iteration 0$\sim$15K) or optimization (Iteration 15$\sim$30K), while the Caching Group remains inactive and does not participate in any calculations.
2 Related Work
In this section, we review relevant work in novel view synthesis and 3D Gaussian Splatting, followed by the distinction of our approach.
$\bullet$ NVS. NVS involves generating new views of a scene from different viewpoints than those in the original dataset. Neural Radiance Fields (NeRF) [20] have been pivotal in NVS for producing photo-realistic images through volume rendering. Neural Radiance Fields (NeRF) have significantly advanced NVS by producing photo-realistic images via volume rendering, but they require substantial computational resources and lengthy training due to costly MLP evaluations. To overcome these limitations, methods such as voxel grids [7], hash grids [22] and points [28] have been introduced to facilitate real-time NeRF rendering.
Recently, 3D Gaussian Splatting (3DGS) [12] has emerged as a promising alternative in the NVS field. Unlike the implicit structure of NeRF, 3DGS uses rasterization rather than ray tracing, enabling efficient rendering suitable for applications requiring both speed and image quality.
$\bullet$ 3DGS. 3D Gaussian Splatting (3DGS) is a recent approach to high-quality, real-time novel view synthesis that represents scenes explicitly as 3D Gaussian primitives. By rasterizing these Gaussians into 2D splats and blending them using $\alpha$-blending [20], 3DGS produces detailed, high-resolution outputs, making it well-suited for applications that demand both speed and visual fidelity.
The 3DGS training process involves two main stages: “Densification” and “Optimization”. During the densification phase (0$\sim$15K iterations), Adaptive Density Control [12] is applied to increase the density of Gaussian primitives, thereby enhancing the model’s capacity for detailed scene representation. In the optimization phase (15Ksimilar-$\sim$30K iterations), densification stops, and the existing Gaussians undergo global refinement, yielding a more precise scene depiction. Recent studies have driven significant advancements in 3DGS across several areas, notably improving rendering quality [2, 30], memory efficiency [24, 5], and scalability for large scenes [15]. Beyond these enhancements, additional research has focused on increasing the efficiency of 3DGS optimizers [10] and refining Gaussian Densification [18, 17] and Pruning strategies [5, 6, 25]. These improvements yield more compact 3DGS models that support faster reconstruction without sacrificing detail.
$\bullet$ Distinction. Unlike previous optimization-focused approaches, we propose Group Training, a principled strategy that dynamically manages Gaussian primitives during training through controlled group sampling. This approach not only accelerates training but also improves reconstruction quality by maintaining effective primitive optimization, achieving a better balance between computational efficiency and visual fidelity.
3 Methodology
In Sec. 3.1, we will provide a comprehensive introduction to the Group Training method and its multifunctional advantages in 3D Gaussian Splatting (3DGS). Section 3.2 introduces two sampling paradigms: Random Sampling and Prioritized Sampling (incorporating importance score, opacity-based, and volume-based criteria), with comparative analysis of their computational efficiency and performance trade-offs. We mathematically demonstrate that Opacity-based Prioritized Sampling achieves effective 3DGS densification and efficient rendering. Finally, in Sec. 3.3 , we summary training characteristics of 3DGS with Group Training, and the Sampling Strategies are identified.
Figure 4: Schedule for activating Group Training during reconstruction. Group Training is enabled at regular intervals. Groups Merging (Grouping without resampling) occurring at iteration $I_{\text{D}}$ and $I_{\text{O}}$ before the completion of the densification and optimization processes for global densification and optimization.
3.1 Group Training
The number of Gaussian primitives in the Gaussian set increases rapidly during the densification phase to achieve rich representation capabilities. This implies that the computational time to render a scene from a specific viewpoint will also increase, and this effect will extend to the global optimisation phase. To this end, we employ the Group Training method during training.
For any given set of Gaussian $G$, the Under-training Group $G_{\text{Under-training}}$ is obtained through sampling, while the remaining primitives constitute the Cached Group $G_{\text{Cached}}$.
$G_{\text{Under-training}}=\left\{g_{i}|g_{i}\in G,i\in I,I\subseteq\{1,2,3,...,|G|\}\right\}\vspace{-12pt}$ (1)
$G_{\text{Cached }}=G\backslash G_{\text{Under-training }}$ (2)
Before the next grouping, the Cached Group is merged with the Under-training Group, after which the new $G_{\text{Under-training}}$ and $G_{\text{Cached}}$ groups are resampled. A visual representation of the Group Training process is provided in Fig. 3 This procedure ensures that each Gaussian primitive has an opportunity to contribute to the training process, and effectively reducing the number of Under-training Gaussians while retaining those deemed “important.”
As outlined in Sec. 2, the 3DGS training process consists of two distinct phases: “Densification” and “Optimization”. Group Training is facilitated in both phases and concludes with Groups Merging at each phases, as illstrated in Fig. 4.
3.2 Effective and Efficient Sampling Strategies
Within each groupinsg, a subset of Gaussians is sampled from the merged Gaussians. The simplest sampling strategy is Random Sampling (RS). We propose attribute-driven Prioritized Sampling to optimize 3DGS further.
We evaluated various attributes , including opacity and volume-based approaches [5] and importance scores [25, 6, 32, 8] (see appendix for detailed comparisons). Among these, Opacity-based Prioritized Sampling demonstrated superior performance in both training speed and rendering quality. As mathematically proven in  Sec. 3.2.1 and 3.2.1, opacity intrinsically governs two critical 3DGS properties, both experimentally validated.
3.2.1 Sampling Strategy for Better Densification
For any Gaussian primitive $G_{m}$ projected onto the imaging plane with 2D center coordinates $[x_{\mathrm{m}},y_{\mathrm{m}}]$. The partial derivatives of loss $L$ w.r.t the projected coordinates are $[\frac{\partial L}{\partial x_{\mathrm{m}}},\frac{\partial L}{\partial y_{\mathrm{m}}}]$, which govern the densification criterion by:
$\sqrt{\left(\frac{\partial L}{\partial x_{\mathrm{m}}}\right)^{2}+\left(\frac{\partial L}{\partial y_{\mathrm{m}}}\right)^{2}}>\tau_{\mathrm{grad}},$
where $\tau_{\mathrm{grad}}$ denotes the predefined gradient threshold.
We formulate the first fundamental property of 3DGS:
$\bullet$ Proposition 1. Opacity-based Effective Gaussians Densification: Gaussian primitives with higher opacity serve as the primary contributors to densification of 3DGS.
Proof. Under our assumptions of mutual independence between Gaussian attributes and within-primitive parameter independence, the partial derivatives are computed as:
$\frac{\partial L}{\partial x_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\end{aligned}$
$\frac{\partial L}{\partial y_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}\frac{\partial\Delta y}{\partial y_{\mathrm{m}}},\end{aligned}$
where $o_{m}$ denotes the opacity of $G_{m}$, $\hat{C}$ represents the rendered pixel value, and $\Delta x,\Delta y$ indicate coordinate offsets between pixels and $[x_{\mathrm{m}},y_{\mathrm{m}}]$. Under the parameter mutual independence assumptions, the $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$, $[\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x},\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}]$ and $\frac{\partial L}{\partial\hat{C}}$ remain independent of $o_{m}$. We prove that $\frac{\partial\hat{C}}{\partial\alpha_{m}}$ increases with higher expected value of $o_{m}$ in appendix:
$\mathbb{E}\left[\frac{\partial\hat{C}}{\partial\alpha_{m}}\right]=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\mathbb{E}[o_{m}]\cdot\mathbb{E}[G_{m}^{\mathrm{2D}}]},$
where $c_{0}$ and $\alpha_{0}$ denote the expected color and opacity respectively, and $T_{\mathrm{saturation}}$ controls $\alpha$ saturation [12].
With fixed $L$Eqs. 4, 5 and 6 demonstrate that elevated opacity distributions yield greater gradient magnitudes in $\left[\frac{\partial L}{\partial x_{\mathrm{m}}},\frac{\partial L}{\partial y_{\mathrm{m}}}\right]$, thereby making such Gaussians statistically prioritized for densification. $\Box$
The distribution of opacity.
The distribution of volume.
Figure 5: The distribution of Gaussian attributes. The distribution of all Gaussian attributes (left) and those contributing specifically to densification (right) in the “Bicycle”[1] with 3DGS [12]. Top row: While the opacities are primarily concentrated around 0 and 1, the Gaussians that contribute to densification are predominantly situated around 1. Bottom row: The distribution of Volume. As densification progresses, Gaussians with lower volume become increasingly involved in the densification process.
$\bullet$ Experimental Validation: Our experiments track Gaussian primitives contributing to densification, focusing on opacity and volumes values. Statistical analysis of attribute distributions during densification steps (Fig. 5) reveals that Gaussians with high opacity and small volume are the primary sources implementing Gaussian densification.
Notably, insufficient high-opacity Gaussians increase the photometric loss, exacerbating both under-reconstruction and over-reconstruction [12], thereby forces redundant densification. While small volumes can be employed as a compression metric [5] in the model compression of 3D Gaussian Splatting, they simultaneously drive reasonable densification. Take into account all factors, it is advisable to cache Gaussians with low-opacity for better gaussian densification.
3.2.2 Sampling Strategy for Faster Rendering
During viewpoint-specific image synthesis in 3DGS, the framework implements $\alpha$-blending [20] through recursive application of the compositing equation on a per-pixel.
$C_{p}=\sum_{i=1}^{N}c_{i}\alpha_{i}T_{i},\quad T_{i}=\prod_{j=1}^{i-1}\left(1-\alpha_{j}\right),$
$N=\min\left\{i\in\mathbb{N}^{+},\bigg{|}\prod_{j=1}^{i-1}(1-\alpha_{j})\leq T_{\mathrm{saturation}}\right\},$
where the parameter $N$ denotes the number of Gaussians required to render pixel $p$. As established in 3DGS [12], the determination of $N$ hinges on achieving opacity saturation within a pixel’s accumulated $\alpha$ value (Transmittance $T_{i}<T_{\mathrm{saturation}}$). This critical threshold implementation introduces a fundamental constraint on rendering efficiency: during forward rendering operations, the sequential traversal of Gaussian primitives inherently prevents parallel computation of $T_{i}$ values. Consequently, $N$ emerges as the principal determinant of rendering speed, where any reduction on $N$ directly translates to accelerated renderings.
Based on the preceding analysis, we formulate the second fundamental property of the 3DGS framework:
$\bullet$ Proposition 2. Opacity-based Efficient Rendering Acceleration: Gaussian primitives with higher opacity enable faster rendering through faster achieving $\alpha$ saturation.
Proof. Building upon the premise of mutually independent $\alpha_{i}$ values between Gaussians, we posit that the mathematical expectation $\mathbb{E}[T_{N}]\approx T_{\mathrm{saturation}}$,
$\mathbb{E}[T_{N}]=(1-\mathbb{E}\left[{\alpha_{i}}\right])^{N}=(1-\mathbb{E}[o_{i}]\cdot\mathbb{E}[G_{i}^{\mathrm{2D}}])^{N},$ (9)
where $\alpha_{i}$ combines its intrinsic opacity $o_{i}$ and $G_{i}^{\mathrm{2D}}$ (dependent on view projection and covariance, independent of $o_{i}$). We decouple the expected value of $\alpha_{i}$. $\Box$
$\bullet$ Experimental Validation: We generated a 3DGS model with Gaussian $o_{i}$ values sampled from $\mathcal{N}(\mu_{o},0.1)$, fixing the other parameters. The rendering time as a function of $\mu_{o}$ is shown in Fig. 6. The results reveals that Gaussian primitives with higher-opacity induce faster transmittance saturation, thereby reducing the blending steps $N$ and rendering computational load.
3.3 Analysis and Design Insights
Through a comprehensive testing program and a detailed examination of the densification and rendering formulas in 3DGS, we have identified the acceleration effect of the opacity attribute in the 3DGS training process. Based on these insights, we propose an Opacity-based Prioritised Sampling (OPS) strategy.
The Opacity-based sampling probability $p_{i}$ for selecting an Under-training Gaussian $G_{i}$ is defined as:
$p_{i}=\frac{\alpha_{i}}{\sum_{i=1}^{N}\alpha_{i}}$ (10)
where $\alpha_{i}$ represents the opacity of Gaussian primitive $G_{i}$, and $N$ is the total number of Gaussian primitives.
Refer to captionRefer to caption
Figure 6: High-opacity Gaussians accelerate 3DGS rendering. Left: The rendering time decreases by 40% as $\mu_{o}$ increases. Right: The mechanisms that highers-opacity Gaussians achieve faster $\alpha$ saturation with less number of Gaussians.
Grouping Iterations Mip-NeRF360 [1] Tanks & Temples [13] Deep Blending [9] Blender [20]
PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
3D Gaussian Splatting* 8.69 792 26.72 4.59 434 15.00 7.72 666 23.90 2.78 69 6.06
Group Training w/ RS 0$\sim$15K 9.51 907 26.62 5.08 496 14.60 7.64 644 21.85 2.76 63 5.66
0$\sim$30K 9.48 902 22.58 5.04 495 12.25 7.63 643 18.45 2.76 63 5.26
Group Training w/ OPS 0$\sim$15K 8.57 678 22.53 4.54 384 12.80 6.81 487 19.25 2.65 43 5.01
0$\sim$30K 8.57 679 19.56 4.55 384 10.95 6.80 487 16.85 2.64 43 4.79
Table 1: Comparison of reconstruction efficiency for 3DGS [12]. Group Training significantly enhances reconstruction speed across all four datasets, with the OPS strategy achieving the highest acceleration. This effect is particularly notable in complex scenes. While Group Training with RS produces a more redundant model, the OPS strategy yields more compact models and reduced GPU peak memory usage. * indicates that the model was retrained. PM stands for GPU peak memory allocation, with Size in MB and Time in minutes.
Grouping Iterations Mip-NeRF360 [1] Tanks&Temples [13] Deep Blending [9] Blender [20]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$
3D GS [12] 27.205 0.815 0.2143 23.142 0.841 0.183 29.405 0.903 0.2425 33.33 0.969 0.030
3D GS* 27.445 0.816 0.2155 26.7 23.697 0.849 0.1764 15.0 29.586 0.904 0.2437 23.9 33.772 0.970 0.0306 6.1
Group Training w/ RS 0$\sim$15K 27.621 0.822 0.2074 26.6 23.773 0.851 0.1703 14.6 29.185 0.901 0.2473 21.9 33.959 0.971 0.0284 5.7
0$\sim$30K 27.537 0.821 0.2157 22.6 23.703 0.848 0.1823 12.3 29.417 0.907 0.2481 18.5 33.877 0.971 0.0295 5.3
Group Training w/ OPS 0$\sim$15K 27.582 0.820 0.2103 22.5 23.842 0.851 0.1726 12.8 29.713 0.906 0.2435 19.3 33.889 0.971 0.0292 5.0
0$\sim$30K 27.564 0.820 0.2133 19.6 23.853 0.850 0.1764 11.0 29.746 0.907 0.2450 16.9 33.808 0.970 0.0299 4.8
Table 2: Comparison of reconstruction quality for 3DGS [12]. Group Training simultaneously accelerates the 3DGS scene reconstruction process and improves the accuracy of 3D reconstructions. Notably, Group Training with OPS achieves a notable improvement in reconstruction quality while maintaining the highest reconstruction speed.
Grouping Iterations Tanks & Temples [13] Deep Blending [9] Blender [20]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Time $\downarrow$
Mip-Splatting* 23.749 0.860 0.1562 23.0 29.358 0.903 0.2390 35.1 33.995 0.970 0.0296 8.5
Group Training w/ RS 0$\sim$15K 23.953 0.863 0.1504 24.6 28.929 0.900 0.2431 31.7 34.203 0.971 0.0278 9.3
0$\sim$30K 24.146 0.862 0.1598 20.4 29.392 0.908 0.2420 26.5 34.123 0.972 0.0285 8.3
Group Training w/ OPS 0$\sim$15K 23.958 0.863 0.1519 21.7 29.665 0.907 0.2364 28.1 34.132 0.971 0.0285 8.2
0$\sim$30K 24.156 0.863 0.1559 18.2 29.788 0.908 0.2384 24.0 34.110 0.971 0.0286 7.5
Table 3: Comparison of reconstruction quality for Mip-Splatting [30]. Group Training achieves simultaneous improvements in both speed and reconstruction quality for Mip-Splatting across all tasks. Specifically, Group Training with OPS attains the fastest reconstruction speed across all three datasets and delivers the highest reconstruction quality in the Tanks & Temples and Deep Blending dataset.
Grouping Iterations Outdoor Indoor Average
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ Time $\downarrow$
Mip-Splatting* 24.814 0.748 0.1995 1430 43.3 31.265 0.929 0.1738 448 33.3 27.681 0.828 0.1881 994 38.8
Group Training w/ RS 0$\sim$15K 24.969 0.757 0.1897 2017 45.9 31.266 0.931 0.1697 556 34.9 27.768 0.834 0.1808 1377 41.0
0$\sim$30K 25.216 0.763 0.1969 2017 36.8 31.164 0.929 0.1757 556 30.5 27.860 0.837 0.1875 1377 34.0
Group Training w/ OPS 0$\sim$15K 24.981 0.754 0.1940 1571 40.7 31.417 0.931 0.1725 375 29.1 27.841 0.832 0.1844 1039 35.5
0$\sim$30K 24.984 0.754 0.1967 1571 32.9 31.407 0.930 0.1739 375 26.4 27.839 0.832 0.1865 1039 30.0
Table 4: Quantitative evaluation on the Mip-NeRF360 [1] reconstructed by Mip-Splatting [30]. For the reconstruction of outdoor scenes, Grouping Training with RS yields a larger model size, resulting in the best reconstruction quality. Grouping Training with OPS achieves the fastest reconstruction speed but provides sub-optimal reconstruction quality. For indoor scenes, Grouping Training with OPS continues to offer both the fastest reconstruction speed and the highest reconstruction quality.
Baseline Baseline with Group Training Ground Truth
3DGS Refer to caption Refer to caption Refer to caption
Time: 26.8 min  PSNR: 28.93 dB Time: 18.6 min  PSNR: 29.47 dB Dr. Johnson [9]
3DGS Refer to caption Refer to caption Refer to caption
Time: 35.9 min  PSNR: 27.21 dB Time: 23.2 min  PSNR: 27.59 dB Garden [1]
Mip-Splatting Refer to caption Refer to caption Refer to caption
Time: 19.5 min  PSNR: 23.66 dB Time: 14.7 min  PSNR: 24.28 dB Train [13]
Mip-Splatting Refer to caption Refer to caption Refer to caption
Time: 54.9 min  PSNR: 24.35 dB Time: 37.7 min  PSNR: 25.18 dB Bicycle [1]
Figure 7: Visual Comparison of Grouping Training and Baseline on 3DGS [12] and Mip-Splatting  [30]. The Grouping Training method accelerates reconstruction by approximately 30% while achieving more detailed and accurate scene rendering. Notable improvements can be observed in the picture frame in the “Dr. Johnson” scene and the bucket handle in the “Garden” scene, where finer details are captured. Furthermore, Grouping Training significantly reduces floaters in Mip-Splatting, enhancing overall reconstruction fidelity.
Cyclic Resample Global Densify Global Optimize Tanks &Temples
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ Size $\downarrow$ Time $\downarrow$
3D GS* 23.697 0.849 0.1764 434 15.0
Group Training 23.866 0.850 0.1749 292 11.8
23.769 0.849 0.1765 231 11.0
23.835 0.850 0.1754 485 11.8
23.853 0.850 0.1764 384 11.0
Table 5: Ablation Study of Group Training with OPS on Tanks & Temples [13]. Cyclic Resampling provides the most substantial acceleration for scene reconstruction, while Global Densification further improves efficiency by reducing model size. With Global Densification enabled, Global Optimization enhances the reconstruction quality of the model.
4 Experiments
We undertake an experimental investigation of Group Training within the context of 3D Gaussian Splatting [12] and Mip-Splatting [30]. The experimental settings are detailed in Sec. 4.1. Subsequently, we present the experimental results in Sec. 4.2. Cross-baselines applicability validation across diverse scenarios is further presented in Sec. 4.3. The ablation experiments and their results are comprehensively discussed in Sec. 4.4. Finally, we analyse the results of different sampling strategies in  Sec. 4.5.
4.1 Experimental Settings
Implementation Details. The Group Training method was integrated as a plug-in to facilitate its utilisation in conjunction with 3DGS. We implemented our method within both 3D Gaussian Splatting [12] [^1] https://github.com/graphdeco-inria/gaussian-splatting and Mip-Splatting [30] [^2] https://github.com/autonomousvision/mip-splatting.
Group Training was implemented with a default grouping interval of 500 iterations throughout the training process. We employed both Random Sampling (RS) and Opacity-Based Prioritized Sampling (OPS) strategies with an Under-Training Ratio (UTR) of 0.6. To ensure computational stability, considering the importance of initial Gaussians [4], Group Training was activated at the 500th iteration after densification begins. The global densification ($I_{\text{D}}$) and optimization ($I_{\text{O}}$) were scheduled at 14.5K and 29K iterations, respectively. To manage the increased GPU memory consumption during group merging, we fixed the SH coefficients [26, 24] at iterations $I_{\text{D}}$ and $I_{\text{O}}$. We also conducted experiments applying Group Training solely during the densification phase.
Dataset and Metrics. The efficacy of our method is evaluated on the Mip-NeRF360 [1] [^3] https://jonbarron.info/mipnerf360/, Tanks & Temples [13] [^4] https://www.tanksandtemples.org/, Deep Blending [9] [^5] http://visual.cs.ucl.ac.uk/pubs/deepblending/, and NeRF-Synthetic [20] [^6] https://www.matthewtancik.com/nerf using a single RTX 3090. Subsequently, the model’s reconstruction performance is evaluated using PSNR, SSIM, and LPIPS. Furthermore, the training time, peak memory usage, and model size for each scene are recorded.
4.2 Experiments Results
This section presents the experimental results of integrating the Group-Training method into the 3D Gaussian Splatting [12] and Mip-Splatting [30]. The test outcomes demonstrate that the Group Training with OPS offers a significant reduction in training time compared to the baseline in all scenes, while concurrently enhancing the rendering quality of the novel view, as presented in Fig. 7.
3D Gaussian Splatting. We first tested the effectiveness of Group Training based on 3DGS. The reconstruction efficiency for all scenes is summarized in Tab. 1, and Tab. 2 presents the NVS rendering quality across all scenes.
Across these scenes, Group Training consistently resulted in faster reconstruction, and this speedup was more pronounced in complex scenes. However, Group Training with RS could lead to an overabundance of redundant Gaussian primitives in some reconstruction tasks, which complicates subsequent operations, such as model compression [24, 23, 26]. Additionally, GPU memory usage significantly increased due to this redundancy [16], as shown in Tab. 1. In comparison to Group Training with RS, Group Training with OPS demonstrated a more pronounced improvement in reconstruction speed. For the reconstruction of all three complex scenes, the average speed increase was about 30%. Moreover, OPS greatly reduced the generation of redundant Gaussian primitives, leading to a smaller reconstruction model size than the original 3DGS reconstruction model in all tasks. Across all reconstruction tasks, the model size was reduced by between 10% and 40%, resulting in a more compact model overall.
Mip-Splatting. We tested Mip-Splatting with Group Training, another significant variant of the Gaussian Splatting model. The experimental results are presented in Tab. 3 and Tab. 4. Overall, the performance of Group Training with Mip-Splatting was similar to 3DGS. Group Training with RS consistently improved both reconstruction quality and speed for Mip-Splatting but produced the highest number of Gaussian primitives during the densification phase. Group Training with OPS achieved the fastest reconstruction speed, and its reconstruction quality was nearly optimal. The OPS reduced the proportion of densification and resulted in a smaller model size that was either the smallest among the tested methods or comparable to the baseline.
Counterintuitive Findings. Our experiments reveal two notable findings that challenge conventional understanding:
  1. 1.
    Despite producing larger model sizes, Group Training with RS consistently improves reconstruction speed, indicating that training dynamics rather than model size dominates reconstruction efficiency.
  2. 2.
    Group Training with OPS simultaneously achieves better reconstruction quality with reduced model size. This demonstrates that superior performance stem from its fundamental impact on the optimization process rather than from model size.
4.3 Methodological Applicability
Group-Training framework fundamentally addresses accelerated training and rendering in 3DGS through its dataset-agnostic improvement. Designed as a pluggable training component, the method demonstrates dual efficacy in both 3DGS acceleration model [6] and compression model [5], which is detailed in appendix.
4.4 Ablation Studies
We conducted an ablation study on the Tanks & Temples dataset [13] to evaluate the impact of three key components of the Group Training method. The detailed comparison results are provided in Tab. 5.
The results indicate that cyclic resampling contributes the most to improving training speed. Specifically, during iterations $I_{\text{D}}$$\sim$15K, the Global Densification process adaptively performs cyclic pruning of low-opacity Gaussians within 3DGS, reducing redundant Gaussians, decreasing the model size, and thereby enhancing training efficiency. When Global Densification is enabled, the 3DGS model removes most redundant Gaussians during global densification, resulting in fewer redundant Gaussians in the Caching Group during the Optimization phase. In this scenario, Global Optimization can make full use of all Gaussians, further improving reconstruction performance.
4.5 Discussion
Our experimental results demonstrate that enabling Group Training and utilizing OPS improved both reconstruction speed and quality. To better clarify these improvements, two key questions arise from these findings:
$\bullet$ Why faser? Group Training significantly accelerates reconstruction primarily by reducing the number of Under-training Gaussian primitives, thereby decreasing the rendering burden on the model. This speedup is observed even when the final model size is larger than the baseline. In alignment with our earlier analysis in Sec. 3.2, the OPS also contributes notably to this speedup effect. During the densification phase, Group Training with OPS preserves high-opacity Gaussians within the Under-Training Group, without which could lead to instances of both under-reconstruction and over-reconstruction [12, 29]. These issues can lead to redundant densification. As a result, models reconstructed with Group Training and OPS contain fewer redundant Gaussians, leading to smaller model sizes and faster reconstruction overall. Additionally, OPS prioritizes sampling high-opacity Gaussians, which rapidly drives $\alpha$ saturation, which also discussed in Sec. 3.2. Focusing on these most impactful Gaussians reduces the number of Gaussians needed for splatting, further enhancing training and rendering efficiency.
$\bullet$ Why better? Group Training with OPS prioritizes the retention and optimization of high-opacity Gaussian primitives, ensuring that elements most crucial to scene rendering are refined more effectively. The Gaussian primitives with higher opacity inherently have a greater influence on the final rendered image [5, 26, 18, 21], so emphasizing their optimization leads to significant improvements in overall rendering quality. Besides, cyclic resampling temporarily excluding a randomly selected subset of Gaussian primitives, each remaining primitive is compelled to learn how to independently represent essential scene features leading to a more robust representation of the scene.
In summary, by integrating these strategies, Group Training with OPS simultaneously enhances both reconstruction speed and rendering quality. By focusing on the most impactful primitives and encouraging independent feature representation, Group Training achieves efficient, high-quality scene reconstruction.
5 Conclusion
This paper presents Group Training, an effective strategy that addresses the computational challenges in 3D Gaussian Splatting by dynamically managing Gaussian primitives in organized groups. Through theoretical analysis and extensive experiments, we demonstrate that our method not only accelerates training by up to 30% but also improves rendering quality across diverse scenarios. Our opacity-based sampling strategy shows universal compatibility with existing 3DGS frameworks, achieving consistent improvements without compromising model complexity. The success of Group Training reveals the importance of dynamic primitive management in 3DGS optimization, suggesting promising directions for future research, such as adaptive grouping strategies and dynamic sampling mechanisms for more complex scenes. Our work provides valuable insights for future developments in efficient novel view synthesis.
Appendix A Proof of Property 1: Opacity-based Effective Gaussians Densification
Under the assumptions of mutual independence between Gaussian attributes and intra-primitive parameter independence, the partial derivatives $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$ for any arbitrary Gaussian admit the following derivation:
[…]
$\frac{\partial L}{\partial x_{\mathrm{m}}}&=\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial G^{\mathrm{2D}}_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}}\\ &=\sum\limits_{\mathrm{pixel}}o_{m}\frac{\partial L}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}}\\ &=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x}\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\end{aligned}$
$\frac{\partial L}{\partial y_{\mathrm{m}}}&=o_{m}\sum\limits_{\mathrm{pixel}}\frac{\partial L}{\partial\hat{C}}\frac{\partial\hat{C}}{\partial\alpha_{m}}\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}\frac{\partial\Delta y}{\partial y_{\mathrm{m}}},\end{aligned}$
where $[\frac{\partial\Delta x}{\partial x_{\mathrm{m}}},\frac{\partial\Delta y}{\partial y_{\mathrm{m}}}]$ remain constant parameters determined by the resolution $[W,H]$; $[\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta x},\frac{\partial G^{\mathrm{2D}}_{m}}{\partial\Delta y}]$ derive from the scale, the rotation and the world coordinates of Gaussian primitives (independent of their opacity); and $\frac{\partial L}{\partial\hat{C}}$ represents the loss gradient with respect to the current pixel value.
Given that $\hat{C}$ is formulated as the composite rendering of $N$ Gaussians in Eq. 14, the derivative $\frac{\partial\hat{C}}{\partial\alpha_{m}}$ admits computation via Eq. 15.
$\hat{C}&=\underbrace{\sum\limits_{i=1}^{m-1}\alpha_{i}c_{i}\prod\limits_{j=1}^{i-1}{(1-\alpha_{j})}}_{\mathrm{Before\,Gaussian}\,m}+\underbrace{\alpha_{m}c_{m}\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}_{\mathrm{Gaussian}\,m}\\ &+\underbrace{\sum\limits_{i=m+1}^{N}\overbrace{\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}^{\mathrm{Before\,G\,}m}\alpha_{i}c_{i}(1-\alpha_{m})\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}}_{\mathrm{After\,Gaussian\,}m}\\ &+\underbrace{\prod\limits_{i=1}^{N}(1-\alpha_{i})c_{\mathrm{bg}}}_{\mathrm{background}},\end{aligned}$
$\frac{\partial\hat{C}}{\partial\alpha_{m}}=\overbrace{\prod\limits_{j=1}^{m-1}{(1-\alpha_{j})}}^{\mathrm{Before\,G\,}m}\left[c_{m}-\sum\limits_{i=m+1}^{N}\alpha_{i}c_{i}\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}\right]-\frac{c_{\mathrm{bg}}T_{N}}{1-\alpha_{m}}\end{aligned}$
Subsequently, the mathematical expectation of this derivative is formally established through Eq. 16
$\mathbb{E}\left[\frac{\partial\hat{C}}{\partial\alpha_{m}}\right]&=\overbrace{(1-\alpha_{0})^{m-1}}^{\mathrm{Before\,G\,}m}\left\{c_{0}-c_{0}\alpha_{0}\sum\limits_{i=m+1}^{N}\mathbb{E}\left[\overbrace{\prod\limits_{j=m+1}^{i-1}{(1-\alpha_{j})}}^{\mathrm{After\,G\,}m}\right]\right\}-\frac{c_{\mathrm{bg}}T_{\mathrm{sta.}}}{1-\alpha_{0}}\\ &=(1-\alpha_{0})^{m-1}\left[c_{0}-c_{0}\alpha_{0}\sum\limits_{i=m+1}^{N}(1-\alpha_{0})^{i-m-1}\right]-\frac{c_{\mathrm{bg}}T_{\mathrm{saturation}}}{1-\alpha_{0}}\\ &=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\alpha_{0}}\\ &=\frac{(c_{0}-c_{\mathrm{bg}})T_{\mathrm{saturation}}}{1-\mathbb{E}\left[o_{i}\right]\mathbb{E}\left[G_{i}\right]},\end{aligned}$
Sampling Strategy Tanks&Temples [13]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
3DGS* 23.730 0.8491 0.176 4.6 430 15.3
Group Training Imp. score 23.672 0.8486 0.174 5.8 593 15.7
Vol. 23.718 0.8462 0.182 5.1 493 12.4
Opac. 23.850 0.8500 0.176 4.5 383 11.0
Vol.+Opac. 23.684 0.8475 0.179 4.8 438 11.9
Table 6: Quantitative evaluation of training efficiency on the Tanks&Temples [13] reconstructed by 3DGS [12]. * indicates that we retrain the model. PM stands for GPU peak memory allocation, with Size in MB and Time in minutes. Imp. score = Importance score based, Vol. = Volume-based, Opac. = Opacity-based, Vol.+Opac. = Volume & Opacity-based.
Sampling Strategy Deep Blending [9]
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
3DGS* 29.503 0.9038 0.244 7.8 677 25.2
Group Training Imp. score 29.589 0.9051 0.246 8.5 765 23.2
Vol. 29.448 0.9036 0.251 7.5 623 19.8
Opac. 29.768 0.9067 0.245 6.8 489 17.2
Vol.+Opac. 29.619 0.9048 0.247 7.0 533 19.0
Table 7: Quantitative evaluation of training efficiency on the Deep Blending [9] reconstructed by 3DGS [12]. Group Training with Opacity-based Prioritised Sampling demonstrates the fastest reconstruction speed and superior performance compared to other sampling strategies.
Mip-NeRF360 [1] Tanks&Temples [13] Deep Blending [9] Blender [20]
PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$ PSNR $\uparrow$ Time $\downarrow$ Accel. $\uparrow$
3D-GS  [12] 27.45 26.7 23.70 15.0 29.59 23.9 33.77 6.1
+Group Training 27.56 19.6 27% 23.85 11.0 27% 29.75 16.9 29% 33.81 4.8 21%
Mini-Splatting  [6] 27.27 20.7 23.26 12.6 29.95 17.8 31.60 10.0
+Group Training 27.25 17.9 13% 23.10 9.9 21% 29.85 14.7 17% 31.98 8.4 16%
LightGaussian  [5] 27.06 27.5 23.09 16.1 27.28 25.9 32.95 6.1
+Group Training 27.34 20.5 25% 23.55 11.9 26% 28.50 19.0 27% 33.18 4.6 24%
Table 8: Quantitative comparisons on different baselines and datasets. Group Training with 3DGS achieves faster reconstruction speed than Mini-Splatting across all datasets. Furthermore, Group Training demonstrates consistent acceleration effects on both 3DGS acceleration model (13$\sim$21speedup on Mini-Splatting [6]) and compression model (24$\sim$27speedup on LightGaussian [5]).Accel. = Acceleration Ratio in training time compared to the baseline.
Appendix B Efficiency And Effectiveness For Various Sampling Strategies
We propose various sampling strategies for Group Training, incorporating Prioritized Sampling based on distinct sampling metrics. The sampling probability for each Gaussian primitive $G_{i}$ is defined as follows:
$p_{i}=\frac{\theta_{i}}{\sum_{i=1}^{N}\theta_{i}},\vspace{-2pt}$ (17)
where $\theta_{i}$ represents the sampling metrics (opacity [5], volume [15] or importance score  [^7] Based on code: https://github.com/fatPeter/mini-splatting.git [25, 6, 32]) of Gaussian primitive $G_{i}$, and $N$ is the total number of Gaussian primitives. We also evaluated the metric which both Opacity and Volume are considered simultaneously, referred to as the Volume & Opacity-based method, as applied in [5]. The sampling metric $\theta_{i}$ for Volume & Opacity-based Prioritized Sampling is computed as follows:
$\theta_{i}=\alpha_{i}\cdot V_{i},$ (18)
where $\alpha_{i}$ represents the opacity and $V_{i}$ represents the volume of Gaussian primitive $G_{i}$.
We conducted experiments using 3D Gaussian Splatting (3DGS) on two datasets: Tanks&Temples [13] and Deep Blending [9], both captured with camera-based systems. The comprehensive comparative results are presented in  Tab. 6 and  Tab. 7. Our results demonstrate that Group Training with Opacity-based Prioritized Sampling (OPS) consistently achieves significant improvements in both reconstruction speed and the quality of 3DGS models. Additionally, the reconstructed models exhibit greater compactness, as evidenced by a marked reduction in redundant Gaussian primitives.
However, volume and importance scores are not the most effective sampling metrics, as they fail to differentiate Gaussians that contribute to densification. This deficiency leads to abrupt vacancies in the Gaussian space under high sampling rates, causing the over-reconstruction and under-reconstruction [12]. Consequently, this exacerbates Gaussian densification, introducing redundancy between newly densified Gaussians and those already cached. The detailed analysis is provided in  Sec. 3.2.
Appendix C Methodological Applicability
We perform comparative validation across two distinct 3DGS architectures: an acceleration-optimized model [6] and a compression-focused LightGaussian [5].
Empirical results demonstrate Group-Training’s consistent efficacy across dataset scales, particularly evidenced by reduced temporal overhead in the Blender [20], as shown in Tab. 8. Crucially, our method synergistically integrates with existing acceleration techniques like Mini-Splatting [6], achieving compounded acceleration gains while providing sustained acceleration for compressed models with concurrent fidelity enhancement.
Appendix D Temporal Evolution of Under-Training Gaussian Primitives.
Figure 8: Comparsion of Under-Training Gaussian Primitives. Our Group-Training methodology selectively trains a subset of Gaussian primitives, demonstrating enhanced computational efficiency while mitigating loss of potentially critical points during opacity reset operations.
We visually compare the quantitative differences in under-training Gaussian primitives between Group-Training and 3DGS during scene reconstruction in Fig. 8. 3DGS with Group-Training reduces the training overhead by avoiding full optimization of all Gaussian primitives. Furthermore, during each opacity reset operation, the proposed method retains a higher proportion of geometrically significant primitives compared to baseline. These retained elements, despite their low-opacity values, preserve critical structural information that contributes to scene geometry fidelity.
Appendix E Detailed Experimental Results for All Scenes
We present the reconstruction results for all scenes using Group Training with Random Sampling (RS) and Opacity-based Prioritized Sampling (OPS), evaluated on 3D Gaussian Splatting (3DGS) [12] and Mip-Splatting [30]. The detailed results are provided in  Tabs. 12, 10, 9, 11, 13, 14, 16 and 15.
The experimental results demonstrate that Group Training consistently delivers significant improvements in both reconstruction speed and quality across all tests, with the acceleration effect being particularly pronounced on complex datasets. Notably, Group Training with OPS achieves the fastest reconstruction times while maintaining optimal or near-optimal reconstruction quality.
We compare the effects of enabling RS and OPS during the Gaussian densification phase. The results indicate that Group Training with RS generates a significantly larger number of Gaussian primitives across all scenarios. For example, when reconstructing the “Bicycle” scene using Mip-Splatting, the high density of Gaussian primitives required the use of an NVIDIA A100 GPU for Group Training with RS. In contrast, Group Training with OPS produces sparser Gaussian primitives while delivering comparable or even superior reconstruction quality. Additionally, the reduced number of Gaussian primitives significantly alleviates the burden on peak memory usage.
Grouping Iterations Dr. Johnson Playroom
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
3D Gaussian Splatting [12] 28.766 0.899 0.244 30.044 0.906 0.241
3D Gaussian Splatting* 29.190 0.901 0.2442 9.0 782 26.8 29.981 0.907 0.2431 6.4 549 21.0
Group Training w/ RS 0$\sim$15K 28.383 0.894 0.2517 8.8 733 24.2 29.987 0.908 0.2429 6.5 554 19.5
0$\sim$30K 28.701 0.902 0.2513 8.8 734 20.3 30.133 0.912 0.2448 6.5 552 16.6
Group Training w/ OPS 0$\sim$15K 29.287 0.903 0.2430 8.1 592 21.7 30.138 0.909 0.2439 5.6 382 16.8
0$\sim$30K 29.309 0.904 0.2451 8.1 594 18.6 30.183 0.910 0.2448 5.6 380 15.1
Table 9: Comprehensive quantitative evaluation results on the DeepBlending [9] reconstructed by 3DGS [12]. RS denotes Random Sampling, and OPS denotes Opacity-based Prioritized Sampling.
Grouping Iterations Mip-NeRF360 [1]
bicycle flowers garden stump treehill bonsai counter kitchen room
PSNR 3DGS [12] 25.246 21.520 27.410 26.550 22.490 31.980 28.700 31.317 30.632
3DGS* 25.205 21.484 27.397 26.620 22.514 32.202 28.980 31.222 31.377
Group Training w/ RS 0$\sim$15K 25.228 21.748 27.552 26.854 22.441 32.430 29.121 31.579 31.634
0$\sim$30K 25.217 21.806 27.463 27.095 22.671 31.975 28.850 31.319 31.438
Group Training w/ OPS 0$\sim$15K 25.260 21.751 27.434 26.809 22.402 32.312 29.031 31.539 31.699
0$\sim$30K 25.219 21.741 27.418 26.830 22.522 32.205 28.973 31.425 31.744
SSIM 3DGS [12] 0.771 0.605 0.868 0.775 0.638 0.938 0.905 0.922 0.914
3DGS* 0.765 0.605 0.866 0.773 0.634 0.942 0.908 0.927 0.919
Group Training w/ RS 0$\sim$15K 0.769 0.619 0.872 0.787 0.638 0.946 0.913 0.930 0.923
0$\sim$30K 0.769 0.616 0.871 0.795 0.642 0.942 0.909 0.927 0.920
Group Training w/ OPS 0$\sim$15K 0.770 0.617 0.869 0.785 0.635 0.945 0.911 0.929 0.922
0$\sim$30K 0.768 0.616 0.868 0.786 0.637 0.944 0.909 0.927 0.921
LPIPS 3DGS [12] 0.205 0.336 0.103 0.210 0.317 0.205 0.204 0.129 0.220
3DGS* 0.2103 0.3355 0.1069 0.2149 0.3240 0.2036 0.2001 0.1261 0.2184
Group Training w/ RS 0$\sim$15K 0.2094 0.3246 0.0985 0.1989 0.3182 0.1936 0.1909 0.1216 0.2110
0$\sim$30K 0.2225 0.3324 0.1061 0.2032 0.3281 0.2008 0.2008 0.1280 0.2198
Group Training w/ OPS 0$\sim$15K 0.2074 0.3262 0.1033 0.2044 0.3203 0.1968 0.1954 0.1245 0.2148
0$\sim$30K 0.2125 0.3307 0.1051 0.2075 0.3226 0.1992 0.1987 0.1272 0.2161
PM 3DGS* 11.6 7.7 11.3 9.0 7.5 7.7 6.7 8.4 8.5
Group Training w/ RS 0$\sim$15K 12.1 8.8 12.0 10.9 9.0 8.4 7.1 8.5 8.8
0$\sim$30K 12.0 8.8 12.0 10.9 9.0 8.4 7.2 8.5 8.7
Group Training w/ OPS 0$\sim$15K 11.1 8.1 10.6 9.4 8.2 7.6 6.4 7.6 8.1
0$\sim$30K 11.1 8.0 10.6 9.3 8.4 7.6 6.4 7.6 8.1
Size 3DGS* 1450 858 1391 1163 896 296 284 428 366
Group Training w/ RS 0$\sim$15K 1516 1019 1474 1457 1106 408 348 439 397
0$\sim$30K 1491 1016 1472 1452 1105 407 349 439 390
Group Training w/ OPS 0$\sim$15K 1192 795 1083 1078 899 267 225 277 287
0$\sim$30K 1195 790 1084 1060 921 267 225 275 291
Time 3DGS* 34.1 24.0 35.9 27.2 24.0 20.5 22.8 28.3 23.7
Group Training w/ RS 0$\sim$15K 31.2 24.2 33.2 28.7 25.0 22.1 24.3 27.6 23.3
0$\sim$30K 24.8 20.1 27.3 23.6 20.3 20.6 21.7 24.2 20.6
Group Training w/ OPS 0$\sim$15K 26.9 21.1 27.4 23.5 22.0 18.5 20.4 22.4 20.6
0$\sim$30K 21.8 17.9 23.2 19.4 18.4 17.4 18.7 20.4 18.8
Table 10: Comprehensive quantitative evaluation results on the Mip-NeRF360 [1] reconstructed by 3DGS [12].
Grouping Iterations Train Truck
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
3D Gaussian Splatting [12] 21.097 0.802 0.218 25.187 0.879 0.148
3D Gaussian Splatting* 21.985 0.815 0.2063 3.6 257 12.5 25.409 0.882 0.1464 5.5 610 17.5
Group Training w/ RS 0$\sim$15K 22.064 0.818 0.2031 3.9 278 11.7 25.482 0.885 0.1375 6.3 714 17.5
0$\sim$30K 21.910 0.812 0.2185 3.8 273 10.1 25.495 0.884 0.1460 6.3 716 14.4
Group Training w/ OPS 0$\sim$15K 22.159 0.818 0.2040 3.6 228 10.7 25.524 0.884 0.1411 5.5 539 14.9
0$\sim$30K 22.156 0.816 0.2104 3.6 227 9.3 25.549 0.884 0.1424 5.5 540 12.6
Table 11: Comprehensive quantitative evaluation results on the Tanks&Temples [13] reconstructed by 3DGS [12].
Grouping Iterations Blender [20]
chair drumps ficus hotdog lego materials mic ship
PSNR 3DGS [12] 33.83 26.15 34.87 37.72 35.78 30.00 35.36 30.80
3DGS* 35.581 26.258 35.481 38.004 36.062 30.461 36.649 31.677
Group Training w/ RS 0$\sim$15K 35.736 26.273 35.494 38.242 36.580 30.675 36.842 31.829
0$\sim$30K 35.637 26.224 35.436 38.142 36.441 30.588 36.786 31.765
Group Training w/ OPS 0$\sim$15K 35.688 26.270 35.487 38.145 36.435 30.569 36.719 31.800
0$\sim$30K 35.623 26.227 35.467 38.017 36.335 30.452 36.654 31.692
SSIM 3DGS* 0.988 0.955 0.987 0.985 0.983 0.960 0.993 0.906
Group Training w/ RS 0$\sim$15K 0.988 0.955 0.987 0.986 0.985 0.962 0.993 0.909
0$\sim$30K 0.988 0.956 0.987 0.987 0.985 0.963 0.993 0.910
Group Training w/ OPS 0$\sim$15K 0.988 0.955 0.987 0.986 0.984 0.962 0.993 0.909
0$\sim$30K 0.988 0.955 0.987 0.986 0.984 0.961 0.992 0.909
LPIPS 3DGS* 0.0104 0.0367 0.0118 0.0201 0.0161 0.0370 0.0064 0.1060
Group Training w/ RS 0$\sim$15K 0.0097 0.0355 0.0117 0.0170 0.0131 0.0340 0.0061 0.0998
0$\sim$30K 0.0107 0.0357 0.0118 0.0184 0.0140 0.0351 0.0063 0.1037
Group Training w/ OPS 0$\sim$15K 0.0099 0.0359 0.0118 0.0181 0.0139 0.0356 0.0064 0.1016
0$\sim$30K 0.0102 0.0364 0.0119 0.0189 0.0143 0.0367 0.0065 0.1042
PM 3DGS* 3.1 2.9 2.7 2.6 2.9 2.6 2.6 2.8
Group Training w/ RS 0$\sim$15K 2.9 2.8 2.6 2.7 3.0 2.6 2.6 2.8
0$\sim$30K 2.9 2.8 2.6 2.7 3.0 2.6 2.6 2.8
Group Training w/ OPS 0$\sim$15K 2.8 2.7 2.5 2.6 2.8 2.6 2.6 2.7
0$\sim$30K 2.8 2.7 2.5 2.6 2.8 2.5 2.6 2.7
Size 3DGS* 116 92 63 44 82 39 46 66
Group Training w/ RS 0$\sim$15K 88 78 39 46 97 47 42 69
0$\sim$30K 87 78 39 47 97 47 43 69
Group Training w/ OPS 0$\sim$15K 59 57 27 31 58 32 30 49
0$\sim$30K 61 58 27 31 58 30 30 49
Time 3DGS* 7.4 6.6 5.3 5.9 6.5 4.9 5.3 6.6
Group Training w/ RS 0$\sim$15K 6.0 5.7 4.5 6.0 6.4 5.2 4.8 6.7
0$\sim$30K 5.5 5.3 4.3 5.5 5.8 4.9 4.6 6.2
Group Training w/ OPS 0$\sim$15K 5.3 5.3 4.1 5.2 5.3 4.6 4.5 5.8
0$\sim$30K 5.1 5.0 4.0 5.0 5.0 4.5 4.2 5.5
Table 12: Comprehensive quantitative evaluation results on the Blender [20] reconstructed by 3DGS [12].
Grouping Iterations Dr. Johnson Playroom
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
Mip-Splatting* 28.711 0.898 0.2431 10.5 981 39.3 30.005 0.907 0.2348 7.4 673 30.8
Group Training w/ RS 0$\sim$15K 27.957 0.892 0.2526 10.1 898 34.3 29.901 0.908 0.2335 8.0 749 29.1
0$\sim$30K 28.500 0.902 0.2486 10.1 902 28.2 30.283 0.914 0.2354 8.0 748 24.7
Group Training w/ OPS 0$\sim$15K 29.145 0.903 0.2393 9.4 732 31.0 30.185 0.910 0.2334 6.6 520 25.2
0$\sim$30K 29.271 0.904 0.2407 9.4 734 26.0 30.305 0.911 0.2360 6.6 518 22.0
Table 13: Comprehensive quantitative evaluation results on the DeepBlending [9] reconstructed by Mip-Splatting [30].
Grouping Iterations Mip-NeRF360 [1]
bicycle flowers garden stump treehill bonsai counter kitchen room
PSNR Mip-Splatting* 25.535 21.753 27.603 26.874 22.304 32.301 29.214 31.803 31.740
Group Training w/ RS 0$\sim$15K 25.664 21.960 27.726 27.124 22.371 32.299 29.291 31.719 31.756
0$\sim$30K 25.784 22.257 27.881 27.446 22.711 32.381 29.084 31.510 31.682
Group Training w/ OPS 0$\sim$15K 25.651 21.852 27.825 27.166 22.411 32.745 29.330 31.870 31.721
0$\sim$30K 25.634 21.844 27.858 27.145 22.441 32.653 29.255 31.864 31.856
SSIM Mip-Splatting* 0.792 0.641 0.877 0.790 0.639 0.945 0.914 0.931 0.925
Group Training w/ RS 0$\sim$15K 0.800 0.653 0.881 0.805 0.647 0.949 0.917 0.932 0.926
0$\sim$30K 0.803 0.656 0.882 0.816 0.658 0.947 0.914 0.931 0.925
Group Training w/ OPS 0$\sim$15K 0.796 0.648 0.879 0.803 0.645 0.948 0.915 0.932 0.926
0$\sim$30K 0.796 0.646 0.879 0.804 0.646 0.948 0.915 0.931 0.926
LPIPS Mip-Splatting* 0.1670 0.2727 0.0950 0.1889 0.2740 0.1881 0.1864 0.1194 0.2011
Group Training w/ RS 0$\sim$15K 0.1607 0.2628 0.0895 0.1736 0.2619 0.1823 0.1807 0.1163 0.1995
0$\sim$30K 0.1708 0.2748 0.0933 0.1751 0.2704 0.1874 0.1891 0.1213 0.2050
Group Training w/ OPS 0$\sim$15K 0.1655 0.2657 0.0922 0.1782 0.2683 0.1843 0.1847 0.1185 0.2026
0$\sim$30K 0.1673 0.2710 0.0929 0.1799 0.2723 0.1861 0.1864 0.1200 0.2029
PM Mip-Splatting* 15.6 9.5 12.3 10.9 10.1 8.4 7.4 9.2 9.7
Group Training w/ RS 0$\sim$15K 19.2 12.6 17.2 15.5 13.7 9.6 8.3 9.9 10.0
0$\sim$30K 19.2 12.7 17.1 15.6 13.7 9.6 8.3 9.9 10.0
Group Training w/ OPS 0$\sim$15K 15.8 10.6 13.9 12.4 11.4 8.4 7.2 8.4 9.0
0$\sim$30K 15.7 10.6 13.9 12.4 11.5 8.4 7.2 8.4 9.0
Size Mip-Splatting* 1957 1089 1475 1398 1232 388 364 523 517
Group Training w/ RS 0$\sim$15K 2494 1550 2194 2083 1763 560 491 615 558
0$\sim$30K 2489 1570 2187 2094 1764 560 490 610 560
Group Training w/ OPS 0$\sim$15K 1968 1230 1684 1564 1410 379 329 388 403
0$\sim$30K 1961 1230 1684 1562 1429 378 329 394 412
Time Mip-Splatting* 54.9 35.4 49.5 39.3 37.3 28.6 31.9 39.0 33.6
Group Training w/ RS 0$\sim$15K 43.7 39.9 57.7 46.3 42.1 32.2 34.1 40.3 32.8
0$\sim$30K 35.1 32.5 45.3 37.0 34.2 28.8 30.0 34.2 28.9
Group Training w/ OPS 0$\sim$15K 47.9 34.3 47.7 37.4 36.2 27.2 28.7 32.0 28.5
0$\sim$30K 37.7 28.3 38.8 30.0 29.5 24.9 26.1 28.5 25.9
Table 14: Comprehensive quantitative evaluation results on the Mip-NeRF360 [1] reconstructed by Mip-Splatting [30].
Grouping Iterations Train Truck
PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$ PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$ PM $\downarrow$ Size $\downarrow$ Time $\downarrow$
Mip-Splatting* 21.783 0.826 0.1892 4.4 351 19.5 25.714 0.893 0.1232 6.8 767 26.5
Group Training w/ RS 0$\sim$15K 22.004 0.829 0.1861 4.8 405 18.9 25.901 0.896 0.1146 9.2 1123 30.3
0$\sim$30K 22.358 0.829 0.1975 4.7 403 16.3 25.934 0.896 0.1221 9.2 1119 24.4
Group Training w/ OPS 0$\sim$15K 21.994 0.830 0.1859 4.4 346 17.4 25.921 0.896 0.1178 7.5 874 25.9
0$\sim$30K 22.167 0.829 0.1915 4.4 348 14.7 25.991 0.895 0.1205 7.6 876 21.1
Table 15: Comprehensive quantitative evaluation results on the Tanks&Temples [13] reconstructed by Mip-Splatting [30].
Grouping Iterations Blender [20]
chair drumps ficus hotdog lego materials mic ship
PSNR Mip-Splatting* 35.773 26.357 35.890 38.267 36.354 30.645 36.934 31.738
Group Training w/ RS 0$\sim$15K 36.078 26.367 35.930 38.474 36.883 30.839 37.152 31.902
0$\sim$30K 35.910 26.293 35.865 38.376 36.860 30.780 36.968 31.931
Group Training w/ OPS 0$\sim$15K 35.981 26.369 35.912 38.396 36.748 30.756 37.067 31.823
0$\sim$30K 35.958 26.363 35.914 38.326 36.739 30.704 37.049 31.824
SSIM Mip-Splatting* 0.988 0.956 0.988 0.986 0.984 0.961 0.993 0.907
Group Training w/ RS 0$\sim$15K 0.989 0.956 0.988 0.987 0.986 0.963 0.993 0.909
0$\sim$30K 0.989 0.957 0.988 0.987 0.986 0.964 0.993 0.911
Group Training w/ OPS 0$\sim$15K 0.989 0.956 0.988 0.987 0.985 0.963 0.993 0.909
0$\sim$30K 0.989 0.956 0.988 0.987 0.985 0.963 0.993 0.909
LPIPS Mip-Splatting* 0.0109 0.0366 0.0111 0.0186 0.0150 0.0359 0.0062 0.1022
Group Training w/ RS 0$\sim$15K 0.0101 0.0355 0.0110 0.0163 0.0126 0.0330 0.0059 0.0980
0$\sim$30K 0.0110 0.0358 0.0111 0.0173 0.0132 0.0341 0.0060 0.0998
Group Training w/ OPS 0$\sim$15K 0.0105 0.0360 0.0111 0.0169 0.0133 0.0343 0.0060 0.0997
0$\sim$30K 0.0105 0.0359 0.0111 0.0171 0.0134 0.0349 0.0061 0.1001
PM Mip-Splatting* 3.0 3.0 2.7 2.7 2.9 2.6 2.8 2.9
Group Training w/ RS 0$\sim$15K 3.3 3.3 2.9 2.9 3.3 2.8 3.0 3.4
0$\sim$30K 3.3 3.4 2.9 2.9 3.3 2.8 3.0 3.4
Group Training w/ OPS 0$\sim$15K 3.1 3.1 2.7 2.7 3.1 2.7 2.9 3.1
0$\sim$30K 3.1 3.1 2.7 2.7 3.1 2.7 2.9 3.1
Size Mip-Splatting* 90 98 51 51 76 40 64 83
Group Training w/ RS 0$\sim$15K 141 146 79 70 136 71 94 146
0$\sim$30K 141 147 80 71 136 71 94 146
Group Training w/ OPS 0$\sim$15K 109 115 57 55 98 51 72 109
0$\sim$30K 109 115 57 55 99 51 72 109
Time Mip-Splatting* 9.2 8.8 6.2 8.4 8.8 6.2 9.3 10.7
Group Training w/ RS 0$\sim$15K 9.6 9.7 6.9 8.8 9.8 7.3 9.8 12.1
0$\sim$30K 8.5 8.7 6.4 8.1 8.8 6.9 8.5 10.8
Group Training w/ OPS 0$\sim$15K 8.6 8.8 6.2 7.8 8.6 6.6 8.4 10.4
0$\sim$30K 7.8 8.0 6.0 7.3 7.7 6.3 7.6 9.6
Table 16: Comprehensive quantitative evaluation results on the Blender [20] reconstructed by Mip-Splatting [30].